# Radon: a Programming Model and Platform for Computing Continuum Systems

### v1

@@ I Introduction
Cloud-computing consolidated as the de-facto standard for building large-scale distributed applications [1]. In this approach, applications are developed as a composition of loosely-coupled services. In fact, managed environments offer application-agnostic platforms with advanced deployment and execution models, such as serverless functions [2], managed databases [3], or data analytics platforms [4, 5].
@@
    Providing an initial assessment of the effectiveness and efficiency of Radon through the implementation of a concrete application – a distributed key-value store
The remainder of the paper is structured as follows. Section II outlines our vision for designing and building Radon. Section III describes Radon’s programming and execution model. Section IV showcases the application of Radon in implementing a distributed key-value store, and evaluates its performance. Finally, Section V reviews related work, and Section VI provides a final overview of the paper, highlighting open research directions.
@@ II Vision
Portability. The Radon runtime offers a common substrate for atoms, making them portable across highly heterogeneous environments. Indeed, the lightweight nature of the runtime makes it suitable to run on diverse hardware architectures, from resource-constrained ones typically found at the edge of the network, to powerful cloud computing infrastructures.
@@
III The Radon Model
Figure 1 shows the high-level architecture of an application developed in Radon. The Radon programming model promotes modularity building on the core concept of atoms. Developers implement their application writing atom definitions, which serve as blueprints and define the behavior of atoms. An atom configuration is then associated with each definition, determining scheduling and deployment choices. By combining an atom definition with its configuration, Radon creates and deploys atom instances across heterogeneous edge-to-cloud environments. Although these instances share the same underlying code (i.e., the definition), each one runs independently in its own memory space and executes concurrently with the others.
Atom instances leverage the services offered by the Radon runtime, which, in analogy with an operating system, plays the role of the kernel that exposes a runtime interface to atoms [^2] To enhance readability and avoid redundancy, the term atom will be used to refer to the general concept. When necessary, explicit distinctions between atom definition and atom instance will be made.. Through this interface, atoms may access a set of runtime functions to control atom execution, communicate with each other (naming and messaging), access storage, set recovery policies, and access additional OS-level features.
Atoms may also internally rely on libraries that extend and enrich the runtime interface. Libraries can be used to abstract features implemented on top of the runtime, which can then be reused across multiple atoms. For instance, a library may implement causally-ordered delivery of messages on top of the FIFO-ordered messaging primitive exposed by the runtime.
In the remainder, we detail the most important functionalities offered by the runtime, such as atom execution (Section III-A), naming and messaging (Section III-B), storage (Section III-C), and recovery (Section III-D). We omit the discussion on other functionalities, such as querying the system’s real-time clock or generating secure random numbers, as they don’t significantly contribute to the relevance of the model.
@@ III-A Atom execution
  • on-demand: a new instance is created for each event
  • on-demand-expire: a new instance is created only if none is available, with instances being reused until an expiration criteria is met (time or number of events handled).
@@ IV-A A distributed and replicated key-value store in Radon
The implementation of the our key-value store is inspired by Anna [7] and its design focuses on scalability and high availability. The system is organized in three atom definitions: KVNodeD, CoordinatorD, and KVFrontend.
KVNodeD sits at the core of the application. It controls storage access and provides a put-get message-based API for key-value pairs. Multiple KVNodeD instances are organized in a ring topology using consistent hashing. Each instance is responsible for one or more key partitions and replicates data across the ring based on a configurable replication factor $N$. Each key-value pair is replicated across N consecutive elements on the ring.
CoordinatorD manages the ring topology by tracking active KVNodeD instances. A single CoordinatorD instance is deployed per application. It responds to topology queries from other atoms and handles additions of ring elements by notifying neighbors to rebalance key partitions.
KVFrontend is the gateway of the system. It process incoming REST requests for put and get operations in the form of events, and it routes them to the correct KVNodeD instance based on key hashes.
CoordinatorD and KVNodeD are both daemon atoms. The former is instantiated on a single node, while the latter is instantiated multiple times on different hosts. KVFrontend is instead a reactive atom instantiated dynamically to serve incoming requests from clients. It has an on-demand-expire policy with an expiration time of 5s.
At startup, each KVNodeD instance contacts the CoordinatorD instance to register itself and obtain the current ring topology. The CoordinatorD instance updates the topology and notifies all existing KVNodeD instances of the changes, prompting key rebalancing. KVFrontend instances fetch the ring topology from CoordinatorD when instantiated.
Our implementation prioritizes high availability over strong consistency. KVFrontend instances send their request to the KVNodeD responsible for the key to be read or updated. If, due to topology changes, the key is not found at that node, it is responsibility of the receiving KVNodeD to forward the request along the ring, until the first responsible node is found. If the request was a put, the KVNodeD updates its local storage and forwards the request to the following KVNodeD instances until all replicas are updated. If the request was a get, the KVNodeD retrieves the value from its local storage and forwards it back to the requesting KVFrontend.
@@ IV-B Performance assessment
We conduct a performance assessment by deploying multiple nodes of our key-value store and simulating concurrent client requests. We measure performance in terms of throughput and latency. We run the experiments on a Workstation equipped with an AMD Ryzen 9 9950X, 64GB of DDR5 RAM, and a 1TB NVMe PCI Gen3$\times$4 DRAM-less QLC SSD, running Fedora Server 41 (kernel 6.13.5). For consistency, we disable Turbo Boost, limiting the CPU clock to 4.3GHz. We compile all Rust code in release mode with rustc 1.85.0 (stable). The Radon runtime uses Wasmer 5.0.4 with LLVM 18.1.8 to compile the atoms using an AoT compilation strategy.
We simulate a distributed scenario by running three instances of the Radon runtime on different ports (we will refer to these three instances as nodes). Each node hosts the components of the key-value store, specifically 8 KVNodeD instances and the on-demand KVFrontend instances. Additionally, one of the nodes runs a single CoordinatorD. A multi-threaded client submits requests to all three nodes. We emulate multiple clients using asynchronous tasks that send requests sequentially, yielding to the asynchronous runtime while waiting for network operations. Each emulated client randomly selects one of the three nodes and directs all its requests to it. Requests are distributed as follows: 20% put, 40% get (random key), 40% get (recently inserted key). Each client limits its requests to reach a target rate per client, if the rate cannot be reached it sends as many requests as it can.
@@
We measure the maximum throughput varying the number of clients (1–128) and the rate of requests from each client (100–10k). Each test runs for 10s. We report the measurements for a replication factor of 2: upon insertion of an item, that item will be stored in 2 KVNodeD instances before the response is sent back to the client. Fig. 2 shows the results we measure: our key-value store can sustain up to about 280k requests/s served across all clients. Further increasing the input rate does not lead to a higher throughput, as requests get queued at the client as it waits for previous responses. In comparison, the baseline radon-echo server reaches about 540k requests/s, while the echo server achieves approximately 700k requests/s. This proves the efficiency of our store: despite the additional forwarding of requests across KVNodeD instances and storage access, the achieved throughput is only a factor of 2 lower than that of a Radon-based server solely echoing HTTP request bodies. Furthermore, the overhead introduced by Radon compared to the vanilla Rust implementation corresponds to a factor of approximately 1.3. This includes the overhead coming from the fact that atoms are compiled to WASM and executed with Wasmer.
@@
Figure 3: Latency distribution across requests for key-value store, radon-echo, and echo.
@@ VI Conclusion and Future Works
Looking ahead, there are several directions for future research: (i) We plan to extend the runtime with asynchronous semantics, addressing the current lack of concurrency at the level of individual atom instances. To support more complex interaction patterns among atoms while preserving their single-threaded execution model, we aim to implement a form of implicit cooperation between atoms, similar to interrupt handling in operating systems. (ii) We intend to broaden Radon’s interaction model beyond REST-based request-response patterns. This includes enabling Radon to proactively interact with the external world, not merely as a reaction to incoming events. (iii) We seek to explore methods and techniques to simplify and automate the deployment and dynamic adaptation of distributed applications, particularly in large-scale and heterogeneous environments. (iv) We aim to build a catalog of components on top of Radon to assess the generality of the proposed API. (v) We aim at assessing the performance of the platform in different computing continuum scenarios.
By addressing these challenges, we believe Radon has the potential to become a tool for building next-generation distributed applications, empowering developers with greater flexibility and efficiency across cloud and edge environments.

### v2

@@ I Introduction
Cloud-computing consolidated as the de-facto standard for building large-scale distributed applications [1]. In this approach, applications are developed as a composition of loosely-coupled services. In fact, managed environments offer application-agnostic platforms with advanced deployment and execution models, such as serverless functions [2], managed databases [3], or data analytics platforms [4, 5, 6].
@@
    Providing an initial assessment of the effectiveness and efficiency of Radon through the implementation of a concrete application – a distributed key-value store.
The remainder of the paper is structured as follows. Sec. II outlines our vision for designing and building Radon. Sec. III describes Radon’s programming and execution model. Sec. IV showcases the application of Radon in implementing a distributed key-value store, and evaluates its performance. Finally, Sec. V reviews related work, and Sec. VI provides a final overview of the paper, highlighting open research directions.
@@ II Vision
Portability. The Radon runtime offers a common substrate for atoms, making them portable across heterogeneous environments. Indeed, the lightweight nature of the runtime makes it suitable to run on diverse hardware architectures, from resource-constrained ones typically found at the edge of the network, to powerful cloud computing infrastructures.
@@ III The Radon Model
Fig. 1 shows the high-level architecture of an application developed in Radon. The Radon programming model promotes modularity through the core concept of atoms. Developers implement their application writing atom definitions, which serve as blueprints and define the behavior of atoms. An atom configuration is then associated with each definition, determining scheduling and deployment choices. By combining an atom definition with its configuration, Radon creates and deploys atom instances across heterogeneous edge-to-cloud environments. Although instances share the same underlying code (i.e., the definition), each one runs independently in its own memory space and executes concurrently with the others.
Atom instances leverage the services offered by the Radon runtime, which, in analogy with an operating system, plays the role of the kernel that exposes a runtime interface to atoms [^2] To enhance readability, we will use the term atom to refer to the general concept, making explicit distinctions between atom definition and atom instance when necessary.. Through this interface, atoms may access a set of runtime functions to control atom execution, communicate with each other (naming and messaging), access storage, set recovery policies, and access additional OS-level features.
Atoms may also internally rely on libraries that build on top of the runtime to extend and enrich its interface, and can then be reused across multiple atoms. For instance, a library may implement causally-ordered delivery of messages on top of the FIFO-ordered messaging primitive exposed by the runtime.
In the remainder, we detail the most important functionalities offered by the runtime, such as atom execution (Sec. III-A), naming and messaging (Sec. III-B), storage (Sec. III-C), and recovery (Sec. III-D). We omit the discussion on other functionalities, such as querying the system’s real-time clock or generating secure random numbers, as they do not significantly contribute to the relevance of the model.
@@ III-A Atom execution
  • on-demand: a new instance is created for each event;
  • on-demand-expire: a new instance is created only if none is available, with instances being reused until an expiration criterion is met (time or number of events handled).
@@ IV-A A distributed and replicated key-value store in Radon
The implementation of the key-value store is inspired by Anna [8] and its design focuses on scalability and high availability. The system is organized in three atom definitions: KVNodeD, CoordinatorD, and KVFrontend.
KVNodeD sits at the core of the application. It controls storage access and provides a put-get message-based API for key-value pairs. Multiple KVNodeD instances are organized in a ring topology using consistent hashing. Each instance is responsible for one or more key partitions and replicates data across the ring based on a configurable replication factor $N$. Each key-value pair is replicated across $N$ consecutive elements on the ring. CoordinatorD manages the ring topology by tracking active KVNodeD instances. A single CoordinatorD instance is deployed per application. It responds to topology queries from other atoms and handles additions of ring elements by notifying neighbors to rebalance key partitions. KVFrontend is the gateway of the system. It processes incoming REST requests for put and get operations in the form of events, and it routes them to the correct KVNodeD instance based on key hashes. CoordinatorD and KVNodeD are both daemon atoms. The former is instantiated on a single node, while the latter is instantiated multiple times on different hosts. KVFrontend is instead a reactive atom instantiated dynamically to serve incoming requests from clients. It has an on-demand-expire policy with an expiration time of 5s. At startup, each KVNodeD instance contacts the CoordinatorD instance to register itself and obtain the current ring topology. The CoordinatorD instance updates the topology and notifies all existing KVNodeD instances of the changes, prompting key rebalancing. KVFrontend instances fetch the ring topology from CoordinatorD when instantiated.
Our implementation prioritizes high availability over strong consistency. KVFrontend instances send their request to the KVNodeD responsible for the key to be read or updated. If, due to topology changes, the key is not found at that node, it is responsibility of the receiving KVNodeD to forward the request along the ring, until the first responsible node is found. If the request is a put, the KVNodeD updates its local storage and forwards the request to the following KVNodeD instances until all replicas are updated. If the request is a get, the KVNodeD retrieves the value from its local storage and forwards it back to the requesting KVFrontend.
@@ IV-B Performance assessment
To conduct a performance assessment, we deployed multiple nodes of our key-value store and simulated concurrent client requests, and we measured throughput and latency of requests. We run the experiments on a Workstation equipped with an AMD Ryzen 9 9950X, 64GB of DDR5 RAM, and a 1TB NVMe PCI Gen3$\times$4 DRAM-less QLC SSD, running Fedora Server 41 (kernel 6.13.5). For consistency, we disable Turbo Boost, limiting the CPU clock to 4.3GHz. We compile all Rust code in release mode with rustc 1.85.0 (stable). The Radon runtime uses Wasmer 5.0.4 with LLVM 18.1.8 to compile the atoms using an AoT compilation strategy.
We simulate a distributed scenario by running three instances of the Radon runtime on different ports (we will refer to these three instances as nodes). Each node hosts the components of the key-value store, specifically 8 KVNodeD instances and the on-demand KVFrontend instances. Additionally, one of the nodes runs a single CoordinatorD. A multithreaded client submits requests to all three nodes. We emulate multiple clients using asynchronous tasks that send requests sequentially, yielding to the asynchronous runtime while waiting for network operations. Each emulated client randomly selects one of the three nodes and directs all its requests to it. Requests are distributed as follows: 20% put, 40% get (random key), 40% get (recently inserted key). Each client limits its requests to reach a target rate per client, if the rate cannot be reached it sends as many requests as it can.
@@
We measure the maximum throughput varying the number of clients (1–128) and the rate of requests from each client (100–10k). Each test runs for 10s. We report the measurements for a replication factor of 2: upon insertion of an item, that item will be stored in 2 KVNodeD instances before the response is sent back to the client. Fig. 2 shows the results we measure: our key-value store can sustain up to about 280k requests/s served across all clients. Further increasing the input rate does not lead to a higher throughput, as requests get queued at the client as it waits for previous responses. In comparison, the baseline radon-echo server reaches about 540k requests/s, while the echo server achieves approximately 700k requests/s. This proves the efficiency of our store: despite the additional forwarding of requests across KVNodeD instances and storage access, the achieved throughput is only a factor of 2 lower than that of a Radon-based server solely echoing HTTP request bodies. Furthermore, the overhead introduced by Radon compared to the vanilla Rust implementation corresponds to a factor of approximately 1.3. This includes the overhead of compiling atoms to WASM and executing the with Wasmer.
@@
Figure 3: Latency distribution for key-value store, radon-echo, and echo.
@@ VI Conclusion and Future Works
Looking ahead, there are several directions for future research: (i) To address the lack of concurrency at the level of individual atom instances while preserving their single-threaded execution model, we plan to implement a form of asynchronous semantics with implicit cooperation between atoms, similar to interrupt handling in operating systems. (ii) We intend to broaden Radon’s interaction model beyond REST-based request-response patterns. This includes enabling Radon to proactively interact with the external world, not merely as a reaction to incoming events. (iii) We will explore methods and techniques to simplify and automate the deployment and dynamic adaptation of distributed applications, particularly in large-scale and heterogeneous environments. (iv) To assess the generality and performance of Radon, we plan to build a catalog of components and by testing their behavior in different computing continuum scenarios.
We believe Radon has the potential to become a tool for building next-generation distributed applications, empowering developers with greater flexibility and efficiency across cloud and edge environments.
Acknowledgements
We acknowledge financial support from the PNRR MUR project PE0000021-E63C22002160007-NEST.